OpenStudy (anonymous):

Solve the system of equations. 4x - y = 8 -8x + 2y = -5

OpenStudy (anonymous):

Multiply the first equation by 2, then add to the 2nd equation. Makes sense?

OpenStudy (anonymous):

so 8x-8y=64?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

by multilyin first eqn by 2 ull get 8x-2y=16 now add both eqn..

OpenStudy (anonymous):

-16x+4y=-10

OpenStudy (anonymous):

Firstly multiply first by 2 you will get: 8x-2y=16 -8x + 2y = -5 Just add them: x and y both cancels.. So, it is the system of linear equation having no solutions because they are parallel to each other and are non - overlapping.. So, this will have no solution..

OpenStudy (anonymous):

yes @waterineyes is correct

OpenStudy (anonymous):

Thank you @arch1

OpenStudy (anonymous):

so 11?

OpenStudy (anonymous):

16+ -5?

OpenStudy (anonymous):

I am saying that the answer to this problem is NO SOLUTION..

OpenStudy (anonymous):

You have to just write like this: \(\large \color{green}{The \:given \:system \:has \;No \:Solution..}\)